# Efficient State Estimation of a Networked FlipIt Model

### v1

@@ Abstract
The Boolean Kalman Filter and associated Boolean Dynamical System Theory have been proposed to study the spread of infection on computer networks. Such models feature a network where attacks propagate through, an intrusion detection system that provides noisy signals of the true state of the network, and the capability of the defender to clean a subset of computers at any time. The Boolean Kalman Filter has been used to solve the optimal estimation problem, by estimating the hidden true state given the attack-defense dynamics and noisy observations. However, this algorithm is infeasible because it runs in exponential time and space with respect to the network size. We address this feasibility problem by proposing a mean-field estimation approach, which is inspired by the epidemic modeling literature. Although our approach is heuristic, we prove that our estimator exactly matches the optimal estimator in certain non-trivial cases. We conclude by using simulations to show both the run-time improvement and estimation accuracy of our approach.
@@ I Introduction
Cybersecurity has been extensively studied with mathematical treatments while assuming full and accurate information (see, e.g. [1, 2]). However, the StuxNet incident [3] served as a “wake-up call” on studying partial and noisy information, dubbed low-information, where partial means defenders’ sensors are only employed at limited places and noisy means the sensors provide inaccurate information. This can be attributed to both the inadequacy of sensor employments (mandated by cost constraints) and the sophistication of attacks, especially the new or zero-da attacks waged by nation-state actors, also known as Advanced Persistent Threats (APTs).
@@
Although many extensions to FlipIt have been proposed, including multiple computers [5] and insider threats [6]. This family of models has two weaknesses. (i) They consider an individual computer, rather than networks. In the real world, computers are networked, resulting in much more complicated scenarios as evidenced by cascading effects and lateral movements (after penetrating into a network). (ii) The low-information assumption does not reflect reality in many settings because Intrusion Detection Systems (IDSes) are widely deployed both on computers (or hosts) and in networks to detect cyber attacks. Although far from perfect, IDSes provide a noisy reading of the true states of computers, which is not modeled by FlipIt-style models.
Recently, the theory of Boolean Dynamical Systems [7] has been applied to study the APT problem with both noisy IDSes and network structures. Originally proposed as a discrete-time dynamical system where each node has a state 00 or $1$ and the dynamics are described by arbitrary modulo 2 operations between connected nodes. The network structures and noisy observations introduce the following non-trivial state estimation problem (in addition to the FlipIt-style decision problem): What is the probability that a given computer is compromised (or infected) at a specific point in time given the network structure and the history of IDS signals (i.e., alerts)? Although this estimation problem has been optimally solved by the Boolean Kalman Filter (BKF) [8], it incurs an exponential time and space in terms of the size of the network because it requires computation over every possible combination of network-wide computer states.
In this work, we propose a novel estimator, dubbed the Mean Field Analysis (MFA) estimator, because its equations resemble the MFA dynamics from networked epidemic spreading models [9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. In the proposed MFA method, we keep track of the probability of infection for each node rather than a distribution over all possible states, leading to a polynomial-time estimation algorithm. Our results are summarized as follows:
@@
Thus, we provide a novel estimator that is both theoretically justified and practical for running on large real-world networks. Notably, our proposed method is suitable for real-time computation on large networks with hundreds of nodes, where the state-of-the-art approach BKF takes hours on graphs with just 14 nodes.
@@ II-A Model
We model the spread of cyber attacks in a computer network where the true state of each node is unknown. This can be considered a hidden Markov model with boolean states and given noisy state observations at each discrete time step. We consider networks with $n$ nodes and let $\{x_{t};t=0,1,2,\dots\}$ be a state process where $x_{t}\in\{0,1\}^{n}$ is the state of the process at time $t$. We use $x_{t,l}$ to denote the state of computer $l$ at time $t$ where $x_{t,l}=1$ means the computer is compromised and $x_{t,l}=0$ means the computer is secure. Let $X=[x^{1},x^{2},\ldots,x^{2^{n}}]$ be an $n\times 2^{n}$ matrix where each $x^{i}\in\{0,1\}^{n}$ represents a unique state. Together, all of the columns $x^{i}$ of $X$ include all possible states in $\{0,1\}^{n}$. Occasionally, it is convenient to abuse notations somewhat and treat $X$ as a set, namely $X=\{x^{1},x^{2},\dots,x^{2^{n}}\}$. We denote a distribution over all $2^{n}$ states as $\Pi\in\Delta(X)$, and $P_{t}\in[0,1]^{n}$ a vector of distributions of each node. To avoid confusion, we use consistent symbols for indexing, $t\in\{0,1,2,\ldots\}$ is a time index, $l,k,r\in V$ are node index, and $i,j\in\{1,2,\ldots,2^{n}\}$ are index over $X$ (e.g., $x^{i},x^{j}\in X$). Due to the amount of indexing required, we only use time indexes, when necessary, to describe the evolution of $x_{t},\Pi_{t}$ and $P_{t}$ over time.
@@
$y_{t,l}=\begin{cases}1\mbox{ if }x_{t,l}=1&\mbox{with probability }p\\ 0\mbox{ if }x_{t,l}=1&\mbox{with probability }1-p\\ 0\mbox{ if }x_{t,l}=0&\mbox{with probability }q\\ 1\mbox{ if }x_{t,l}=0&\mbox{with probability }1-q\\ \end{cases}$ (2)
@@ II-B The Boolean Kalman Filter
The optimal estimator on the preceding problem, known as the Boolean Kalman Filter (BKF), has been given in [19] (with minor technical differences). Before giving the full details of the algorithm, we discuss the high-level ideas behind it. The algorithm has three steps at each iteration, where the probability distribution $\Pi_{t-1|t-1}\in\Delta(X)$ is updated to $\Pi_{t|t}\in\Delta(X)$, which is leveraged to produce the optimal MSE estimator. The distribution $\Pi_{t-1|t-1}$ is known as the prior belief and is the existing belief across all states $X$. In the first step, the matrix-vector product is taken
@@
where $M_{t}$ is a $2^{n}\times 2^{n}$ column stochastic matrix that represents the expected dynamics of (1) with the cleaning vector $a_{t-1}$. This produces an intermediary belief $\Pi_{t|t-1}$, which is then operated on by vector $T_{t}(y_{t})$ to obtain the updated belief $\Pi_{t\mid t}$, as follows:
@@
where the overline operator $\overline{\cdot}$ is the element-wise rounding operator. The drawback of the BKF is that all three components, $\Pi_{t\mid t}$,$M_{t}$,$T_{t}$, are all exponentially sized, namely at the magnitude of $2^{n}$. This automatically implies the BKF algorithm is not feasible for sufficiently large networks.
@@ III-A An Approximate Approach
$\begin{cases}\frac{\Pr(y_{t+1,l}=1\mid P_{t},a_{t},x_{t+1}=1)\Pr(x_{t+1,l}=1\mid P_{t},a_{t})}{Pr(y_{t+1,l}=1\mid P_{t},a_{t})}&y_{t+1,l}=1\\ \frac{(1-\Pr(y_{t+1}=1\mid P_{t},a_{t},x_{t+1}=1))\Pr(x_{t+1,l}=1\mid P_{t},a_{t})}{(1-Pr(y_{t+1,l}=1\mid P_{t},a_{t}))}&y_{t+1,l}=0\end{cases}$
via Bayes’ Law. It then suffices to derive each of the three terms as a function of $P_{t},u_{t},y_{t+1}$ to complete our derivation of the MFA estimator. The first term of the numerator is easy to show directly, $\Pr(y_{t+1,l}=1\mid P_{t},a_{t},x_{t+1}=1)=p$; then, by using this result and the law of total probability, the denominator can be solved as
@@ 1 $P^{\prime}_{t,l}\leftarrow(1+(\alpha-1)a_{t,l})$
Given that we cannot directly evaluate Bayes’ Law with belief $P_{t}$, the main premise of our estimator is to leverage the Mean Field Analysis (MFA) equation to provide a tractable equation in place of $\Pr(x_{t+1,l}=1\mid P_{t},a_{t})$. The MFA equations have long been proposed for a similar purpose in epidemic modeling (see, e,g,m [9, 15]), where each node is modeled individually. The dynamics are given to (1), with the main difference that in epidemic modeling, there is no notion of cleaning $a_{t}$, but rather some natural recovery process is assumed. However, both models lead to dynamics similar to (5) where there exists an underlying $2^{n}$ state Markov chain that is difficult to analyze. Thus, the so-called MFA equation has been proposed, given in Line 2 of Algorithm 2. Particularly, it can be seen that the MFA equation is structurally similar to dynamics (1) with the exception that states $x_{t}$ have been replaced with belief $P_{t}$. With respect to networked models, there have been considerable numerical efforts to evaluate the performance of Mean Field estimators [16, 18], and it has been shown via two prominent epidemic models [10, 11] that in a special case, the variations of the MFA dynamics do converge to the true dynamics.
We now provide a high-level description of the proposed Algorithm 2that aims to approximate the results of the BKF Algorithm 1. Both algorithms function in 3 sub-steps (Line 1-3) at each iteration; each step of the MFA algorithm functions analogously to the BKF. Note that the MFA algorithm requires an extra loop over $V$ as the matrix products can no longer be used to do that implicitly. Further, it can be seen that Line 2 utilizes the MFA equation to estimate the expected dynamics (1), given prior belief $P_{t-1}$. This produces an intermediate distribution $P^{\prime}_{t}$, which is then updated in Line 2 with respect to the received observation $y_{t}$ utilizing Bayes’ law as given in (11). This update takes into consideration the proposed approximation $\Pr(x_{t,l}=1\mid P_{t-1},a_{t})\approx P^{\prime}_{t}$. Finally, in Line 2, we round the posterior belief $P_{t}$ in an element-wise fashion to obtain the MFA estimator $\hat{x}_{t}$. It should be noted that, at each time step, the algorithm requires an operation over each node $l\in V$, and within that loop, a loop over all in-neighbors of $l$. Thus, the MFA algorithm can be pessimistically regarded as running in $\mathcal{O}(nD^{*})$ time, where $D^{*}=\max_{l\in V}|D_{l}|$ is the maximum number of in-neighbors among all nodes.
@@ IV Results
Theorem 1 states that the distribution $\Pi^{*}(P)$ that maximizes entropy while being consistent with the belief $P$ is the product distribution over $X$ induced by $P$. The proof of this result is deferred to the Appendix (Section VI). Using this form, we now evaluate $\Pr(x_{t+1,l}=1\mid\Pi^{*}(P_{t}),a_{t})=XM_{t}\Pi^{*}(P_{t})$, which is Line 1 of the BKF Algorithm 1.
@@
Let $M_{t}$ be constructed appropriately via (9) for any $a_{t-1}$. Then, $\Pr(x_{t+1,l}=1\mid\Pi^{*}(P_{t}),a_{t})$ may be computed explicitly in $\mathcal{O}(n)$ via
$\Pr(x_{t+1,l}=1\mid\Pi^{*}(P_{t}),a_{t})=(XM_{t}\Pi^{*}(P_{t}))_{l}$ (14)
@@
The proof is deferred to the Appendix (Section VI). Next, we proceed directly to Line 1, showing that it may also be derived from the BKF, given $\Pi_{t\mid t-1}=\Pi^{*}(P^{\prime}_{t})$.
@@
Given a distribution $\Pi_{t\mid t-1}$ such that $\Pi_{t\mid t-1}=\Pi^{*}(P)$ for some $P\in[0,1]^{n}$, we may compute (6) in $\mathcal{O}(n)$ via
$P_{t}=(X\Pi_{t\mid t})_{r}=\bigg{(}X\frac{T_{t}\circ\Pi_{t\mid t-1}}{(T_{t})^{\top}\Pi_{t\mid t-1}}\bigg{)}_{r}=$ (15)
@@
Again, we defer the proof to the Appendix (Section VI Interestingly, this indicates that the Bayesian update of the observation is equivalent to the BKF’s observation update in this context.
@@ IV-A Simulation Results
We now provide simulations to verify the effectiveness of the MFA algorithm based on the intuition of Theorems 2 and 3. To evaluate the algorithms, we define a True Estimation Rate metric, $\textrm{TER}(x_{t},\hat{x}_{t})=1-(||x_{t}-\hat{x}_{t}||^{2}_{2})/n$, which is the fraction of nodes whose states are correctly estimated. Throughout (unless otherwise stated), the parameters are $\alpha=0.2,p=0.8,q=0.8,\rho_{ij}=0.1$, and we run each experiment for 20 time steps in 100 independent random trials. We assume that the first node is the external threat that is known to be compromised and cannot be cleaned (we also exclude this node for evaluation purposes). The initial state is then $x_{0}=(1,0,0,0,\dots)$; the initial beliefs of the algorithms are $P_{0}=(1,0.5,0.5,\dots)$ and $\Pi_{0}=\Pi^{*}(P_{0})$ for MFA and BKF, respectively. Additionally, at each time step, we assume 2 nodes are randomly selected for cleaning (excluding node 1).

### v2

@@ Abstract
The Boolean Kalman Filter and associated Boolean Dynamical System Theory have been proposed to study the spread of infection on computer networks. Such models feature a network where attacks propagate through, an intrusion detection system that provides noisy signals of the true state of the network, and the capability of the defender to clean a subset of computers at any time. The Boolean Kalman Filter has been used to solve the optimal estimation problem, by estimating the hidden true state given the attack-defense dynamics and noisy observations. However, this algorithm is intractable because it runs in exponential time and space with respect to the network size. We address this feasibility problem by proposing a mean-field estimation approach, which is inspired by the epidemic modeling literature. Although our approach is heuristic, we prove that our estimator exactly matches the optimal estimator in certain non-trivial cases. We conclude by using simulations to show both the run-time improvement and estimation accuracy of our approach.
@@ I Introduction
Cybersecurity has been extensively studied with mathematical treatments while assuming full and accurate information (see, e.g. [1, 2]). However, the StuxNet incident [3] served as a “wake-up call” on studying partial and noisy information, dubbed low-information, where partial means defenders’ sensors are only employed at limited places and noisy means the sensors provide inaccurate information. This can be attributed to both the inadequacy of sensor employments (mandated by cost constraints) and the sophistication of attacks, especially the new or zero-day attacks waged by nation-state actors, also known as Advanced Persistent Threats (APTs).
@@
Although many extensions to FlipIt have been proposed, including multiple computers [5], deployment of honeypots [6], insider threats [7], and even computer networks [8, 9], but there are limitations to the practical application of these models. Primairly, the low-information assumption does not reflect reality in many settings because Intrusion Detection Systems (IDSes) are widely deployed both on computers (or hosts) and in networks to detect cyber attacks. Although far from perfect, IDSes provide a noisy reading of the true states of computers, which is not modeled by FlipIt-style models.
Recently, the theory of Boolean Dynamical Systems [10] has been applied to study the APT problem with both noisy IDSes and network structures. Originally proposed as a discrete-time dynamical system where each node has a state 0 or $1$ and the dynamics are described by arbitrary modulo 2 operations between connected nodes. The network structures and noisy observations introduce the following non-trivial state estimation problem (in addition to the FlipIt-style decision problem): What is the probability that a given computer is compromised (or infected) at a specific point in time given the network structure and the history of IDS signals (i.e., alerts)? Although this estimation problem has been optimally solved by the Boolean Kalman Filter (BKF) [11], it incurs an exponential time and space in terms of the size of the network because it requires computation over every possible combination of network-wide computer states.
In this work, we propose a novel estimator, dubbed the Mean Field Analysis (MFA) estimator, because its equations resemble the MFA dynamics from networked epidemic spreading models [12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. In the proposed MFA method, we keep track of the probability of infection for each node (which requires $n$ real numbers) rather than a distribution over all possible states (which requires $2^{n}$ real numbers), leading to a polynomial-time estimation algorithm. Our results are summarized as follows:
@@
Thus, we provide a novel estimator that is both theoretically justified and practical for running on large real-world networks. Notably, our proposed method is suitable for real-time computation on large networks with hundreds of nodes, where the state-of-the-art approach BKF takes hours on graphs with just 14 nodes. Notably, a previous approach also developed a polynomial-time algorithm to estimate the BKF based on particle filters [22]. In contrast to their approach, our estimation approach leverages specific problem characteristics to draw parallels with MFA as it is applied in epidemic theory.
@@ II-A Model
We model the spread of cyber attacks in a computer network where the true state of each node is unknown. This can be considered a hidden Markov model with boolean states and given noisy state observations at each discrete time step. We consider networks with $n$ nodes and let $\{x_{t};t=0,1,2,\dots\}$ be a state process where $x_{t}\in\{0,1\}^{n}$ is the state of the process at time $t$. We use $x_{t,l}$ to denote the state of computer $l$ at time $t$ where $x_{t,l}=1$ means the computer is compromised and $x_{t,l}=0$ means the computer is secure. Let $X=[x^{1},x^{2},\ldots,x^{2^{n}}]$ be an $n\times 2^{n}$ matrix where each $x^{i}\in\{0,1\}^{n}$ represents a unique state. Together, all of the columns $x^{i}$ of $X$ include all possible states in $\{0,1\}^{n}$. Occasionally, it is convenient to abuse notations somewhat and treat $X$ as a set, namely $X=\{x^{1},x^{2},\dots,x^{2^{n}}\}$. We denote a distribution over all $2^{n}$ states as $\Pi\in\Delta(X)$, where $\Delta(X)=\{\Pi\in[0,1]^{2^{n}}\mid\sum\Pi=1\}$ is the space of all distributions over $X$. Further, $P_{t}\in[0,1]^{n}$ a vector of distributions of each node, where $P_{t},l$ encodes the marginal belief that node $l$ is infected at time $t$. To avoid confusion, we use consistent symbols for indexing, $t\in\{0,1,2,\ldots\}$ is a time index, $l,k,r\in V$ are node indexes, and $i,j\in\{1,2,\ldots,2^{n}\}$ are indexes over $X$ (e.g., $x^{i},x^{j}\in X$). We will use the indexes $i,j$ from $x^{i},x^{j}$ to refer to the related probability in $\Pi$ as $\Pi_{i},\Pi_{j}$. Due to the amount of indexing required, we only use time indexes, when necessary, to describe the evolution of $x_{t},\Pi_{t}$ and $P_{t}$ over time.
@@
[…]
@@ II-B The Boolean Kalman Filter
The optimal estimator on the preceding problem, known as the Boolean Kalman Filter (BKF), has been given in [23] (with minor technical differences). Before giving the full details of the Algorithm 1, we discuss the high-level ideas behind it. The algorithm has three steps at each iteration, where the probability distribution $\Pi_{t-1|t-1}\in\Delta(X)$ is updated to $\Pi_{t|t}\in\Delta(X)$, which is leveraged to produce the optimal MSE estimator. The distribution $\Pi_{t-1|t-1}$ is known as the prior belief and is the existing belief across all states $X$. In the first step, the matrix-vector product is taken
@@
where $M_{t}$ is a $2^{n}\times 2^{n}$ column stochastic matrix that represents the expected dynamics of (1) with the cleaning vector $a_{t-1}$. which is then operated on by observation information (encoded as vector $T_{t}(y_{t})$) to obtain the updated belief $\Pi_{t\mid t}$, as follows:
@@
where the overline operator $\overline{\cdot}$ is the element-wise rounding operator. The drawback of the BKF is that all three components, $\Pi_{t\mid t}$$M_{t}$$T_{t}$, are all exponentially sized, namely at the magnitude of $2^{n}$. This automatically implies the BKF algorithm is not feasible for sufficiently large networks.
@@ III-A An Approximate Approach
[…]
via Bayes’ Law. It then suffices to derive each of the three terms as a function of $P_{t},a_{t},y_{t+1}$ to complete our derivation of the MFA estimator. The first term of the numerator is easy to show directly, $\Pr(y_{t+1,l}=1\mid P_{t},a_{t},x_{t+1}=1)=p$; then, by using this result and the law of total probability, the denominator can be solved as
@@ 1 $P^{\prime}_{t,l}\leftarrow(1+(\alpha-1)a_{t,l})$
Given that we cannot directly evaluate Bayes’ Law with belief $P_{t}$, the main premise of our estimator is to leverage the Mean Field Analysis (MFA) equation to provide a tractable equation in place of $\Pr(x_{t+1,l}=1\mid P_{t},a_{t})$. The MFA equations have long been proposed for a similar purpose in epidemic modeling (see, e.g. [12, 18]), where each node is modeled individually. The dynamics are given to (1), with the main difference that in epidemic modeling, there is no notion of cleaning $a_{t}$, but rather some natural recovery process is assumed. However, both models lead to dynamics similar to (5) where there exists an underlying $2^{n}$ state Markov chain that is difficult to analyze. Thus, the so-called MFA equation has been proposed, given in Line 2 of Algorithm 2. Particularly, it can be seen that the MFA equation is structurally similar to dynamics (1) with the exception that states $x_{t}$ have been replaced with belief $P_{t}$. With respect to networked models, there have been considerable numerical efforts to evaluate the performance of Mean Field estimators [19, 21], and it has been shown via two prominent epidemic models [13, 14] that in a special case, the variations of the MFA dynamics do converge to the true dynamics.
We now provide a high-level description of the proposed Algorithm 2 that aims to approximate the results of the BKF Algorithm 1. Both algorithms function in 3 sub-steps (Line 1-3) at each iteration; each step of the MFA algorithm functions analogously to the BKF. Note that the MFA algorithm requires an extra loop over $V$ as the matrix products can no longer be used to do that implicitly. Further, it can be seen that Line 2 utilizes the MFA equation to estimate the expected dynamics (1), given prior belief $P_{t-1}$. This produces an intermediate distribution $P^{\prime}_{t}$, which is then updated in Line 2 with respect to the received observation $y_{t}$ utilizing Bayes’ law as given in (11), producing the new belief $P_{t}$. Finally, in Line 2, we round the posterior distribution $P_{t}$ in an element-wise fashion to obtain the MFA estimator $\hat{x}_{t}$.
It should be noted that, at each time step, the algorithm requires an operation over each node $l\in V$, and within that loop, a loop over all in-neighbors of $l$. Thus, the MFA algorithm can be pessimistically regarded as running in $\mathcal{O}(nD^{*})$ time, where $D^{*}=\max_{l\in V}|D_{l}|$ is the maximum number of in-neighbors among all nodes. Additionally, it can be seen that our MFA algorithm can be seen as a distributed algorithm in the sense that, for any node $l$ to compute their next estimate $P_{t+1,l}$, they only require parameters relating to their in-neighbors $D_{l}$.
@@ IV Results
Theorem 1 states that the distribution $\Pi^{*}(P)$ that maximizes entropy while being consistent with the belief $P$ is the product distribution over $X$ induced by $P$. Due to space constraints we defer the proof to an online version [25]. Using this form, we now evaluate $\Pr(x_{t+1,l}=1\mid\Pi^{*}(P_{t}),a_{t})=XM_{t+1}\Pi^{*}(P_{t})$, which is Line 1 of the BKF Algorithm 1.
@@
Let $M_{t}$ be constructed appropriately via (9) for any $a_{t-1}$. Then, $\Pr(x_{t+1,l}=1\mid\Pi^{*}(P_{t}),a_{t})$ is computed explicitly in $\mathcal{O}(n)$ via
$\Pr(x_{t+1,l}=1\mid\Pi^{*}(P_{t}),a_{t})=P^{\prime}_{{t+1},l}=(XM_{t+1}\Pi^{*}(P_{t}))_{l}$ (14)
@@
The proof is again deferred to the online version [25]. Next, we proceed directly to Line 1, showing that it may also be derived from the BKF, given $\Pi_{t\mid t-1}=\Pi^{*}(P^{\prime}_{t})$.
@@
Given a distribution $\Pi_{t\mid t-1}$ such that $\Pi_{t\mid t-1}=\Pi^{*}(P)$ for some $P^{\prime}_{t}\in[0,1]^{n}$, then (6) is computed in $\mathcal{O}(n)$ is computed via
$P_{t,l}=(X\Pi_{t\mid t})_{l}=\bigg{(}X\frac{T_{t}\circ\Pi_{t\mid t-1}}{(T_{t})^{\top}\Pi_{t\mid t-1}}\bigg{)}_{l}=$ (15)
@@
Again, we defer the proof to [25]. Interestingly, this indicates that the Bayesian update of the observation is equivalent to the BKF’s observation update in this context.
@@ IV-A Simulation Results
[…]